OpenStudy (anonymous):

Find all real solutions. 2*sinx*cosx+4*sinx-cosx-2=0

OpenStudy (anonymous):

we can factor: 2sin(x)(cos(x)+2)-(cos(x)+2)=0 (2sin(x)-1)(cos(x)+2)=0

OpenStudy (anonymous):

sin(x)=1/2

OpenStudy (anonymous):

x=Pi/6, 5*Pi/6

OpenStudy (anonymous):

cos(x) cannot be -2 so these are your solutions on the interval [0, 2Pi)

OpenStudy (anonymous):

ok, thanks. I didn't know where to begin. thank you for all your help.

OpenStudy (anonymous):

you're welcome